OpenStudy (anonymous):

Write the equation of the line perpendicular to the line x - 3y = 9 and passing through the point (3,5).

OpenStudy (anonymous):

the product of slopes of perpendicular lines =-1 -3y=-x+9 y=(1/3)x-3 so the slope of perpendicular line : -3 y-y1=m(x-x1) y-5=-3(x-3) y-5=-3x+9 y=-3x+14